A 24.0 V battery is connected to a circuit that has an ammeter (A1), a 6.0 ohm

resistor connected in series with a voltmeter (V1) and 3 resistors connected in parallel

R1, 2.0 ohms, R2, 2.0 ohms and R3, 1.0 ohms. There is a voltmeter (V2) placed across

the R1 and an ammeter (A2) in line with R2. A. Calculate the values of the meters. (V1,

V2, A1 & A2).


1
Expert's answer
2021-05-24T16:52:07-0400

Resistance of parallel circuit:

"\frac{1}{R_p}=\frac{1}{R_1}+\frac{1}{R_2}+\frac{1}{R_3}"

"R_p=\frac{R_1R_2R_3}{R_2R_3+R_1R_3+R_1R_2}=\frac{2\cdot2}{2+2+4}=0.5\ \Omega"


Then:

Value of A1:

"I=\frac{24}{6+0.5}=3.7\ A"

Value of V1:

"U=3.7\cdot6=22.2\ V"


Voltage in parallel circuit:

"U_p=24-22.2=1.8\ V"


Then:

Value of V2:

"U_p=1.8\ V"

Value of A2:

"I_2=U_p/R_2=1.8/2=0.9\ \Omega"